The sum of three consecutive multiples of  8 is 888. Find the multiple.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the multiples be x, x+8, x+16

Given, the sum of three consecutive multiples of 8 is 888

Thus,

x+x+8+x+16=888

3x+24=888

3x=888-24

3x =864

x =864/3

x =288

Therefore the multiples are:

x =288

x +8=296

x+16=304


Thus , the multiples are 288,296,304.
Thanks